OpenStudy (anonymous):

Give the values of a, b, and c needed to write the equation's general form. (5 + x)(5 - x) = 7 A = 1; B = 0; C = -18 A = 25; B = 0; C = -1 A = -1; B = 0; C = 25

OpenStudy (anonymous):

(5+x)(5-x) = 7 You know that (a+b)(a-b)=a^2 - b^2 So, 25 - x^2 = 7 (or) -x^2 + 18 = 0 (or) x^2 - 18 = 0 So, A=1, B=0 and C= -18